# CogNav: Cognitive Process Modeling for Object Goal Navigation with LLMs

### v2

@@ 3.1 Cognitive Map Construction
[…]
@@ 3.3 Cognitive Process Modeling
[…]
@@ 4.1 Experiment Setup
Implementation Details. For each episode, we set the maximum navigation step to 500. The observation of the agent is $640\times 480$ RGB-D images,, with depth values ranging from $0.5$m to $5$m. The camera of agent is $0.9$m above the ground, with each forward action advancing $0.25$m and each rotation covering $30^{\circ}$. The occupancy map is configured as $960\times 960$ with a resolution of $0.05$m. For instance detection and segmentation, we utilize OpenSEED [44] as our 2D foundation model. In the cognitive process, GPT-3 [3] is used as the large language model (LLM), and GPT-4v [1] serves as the vision-language model (VLM) for decision-making during navigation.
@@ 4.4 Ablation Study
[…]
@@
[…]

### v3

@@ 3.1 Cognitive Map Construction
Occupancy map. To provide spatial scene information, such as explored areas, unexplored areas, and frontiers, we construct a top-down occupancy map $\mathcal{M}_{t}\in\mathbb{R}^{M\times M\times 4}$ from $I_{t}$, following the approach in [7]. $M\times M$ defines the map size with a grid resolution of $5\text{cm}$. The first three channels indicate occupied areas, explored regions, and the agent location, respectively. The fourth channel incorporates the projected 3D instance nodes $\mathcal{N}_{t}$. The occupancy map $\mathcal{M}_{t}$ is updated at each navigation step.
@@ 3.3 Cognitive Process Modeling
SR $\left(\%\right)\uparrow$ SPL $\left(\%\right)\uparrow$ SR $\left(\%\right)\uparrow$ SPL $\left(\%\right)\uparrow$ SR $\left(\%\right)\uparrow$ SPL $\left(\%\right)\uparrow$
@@ 4.1 Experiment Setup
Implementation Details. For each episode, we set the maximum navigation step to 500. The observation of the agent is $640\times 480$ RGB-D images,, with depth values ranging from 0.50.5m to 55m. The camera of agent is 0.90.9m above the ground, with each forward action advancing 0.250.25m and each rotation covering $30^{\circ}$. The occupancy map is configured as $960\times 960$ with a resolution of 0.050.05m. For instance detection and segmentation, we utilize OpenSEED [44] as our 2D foundation model. In the cognitive process, GPT-3 [3] is used as the large language model (LLM), and GPT-4v [1] serves as the vision-language model (VLM) for decision-making during navigation.
@@ 4.4 Ablation Study
SR $\left(\%\right)\uparrow$ SPL $\left(\%\right)\uparrow$ DTG $\left(m\right)\downarrow$
@@
SR $\left(\%\right)\uparrow$ SPL $\left(\%\right)\uparrow$ DTG $\left(m\right)\downarrow$
@@ 5 Conclusion and Future Work
6 Acknowledgements
This work was supported in part by the NSFC (62325211, 62132021, 62402516), the Major Program of Xiangjiang Laboratory (23XJ01009), Key R&D Program of Wuhan (2024060702030143).